# Derivative-Aligned Anticipation of Forbush Decreases from Entropy and Fractal Markers

### v1

@@ Abstract
We present a feature–based framework to anticipate Forbush decreases (FDs) from one–minute neutron–monitor data by tracking sliding–window invariants grounded in information theory, scaling, and geometry. For each station and event, we compute marker time series (Shannon, spectral, permutation, sample and approximate entropy; Lempel–Ziv complexity; Hurst via detrended fluctuation analysis; correlation dimension; Higuchi, Katz, and Petrosian fractal dimensions), smooth them with an exponentially weighted moving average, and analyze their within–station standardized derivatives relative to the time of the minimum of the smoothed count derivative. Leads are reported in minutes (negative = precedes) and station–level significance is operationally defined on a pre–onset window by a robust zz–score detector with bilateral threshold and persistence; no cross–correlation or hypothesis testing is required. Applied to three recent FDs (2023-04-23, 2024-03-24, 2024-05-10), the analysis reveals a compact, robust panel—Hurst (DFA), Katz fractal dimension, Shannon and sample entropy, and Lempel–Ziv complexity—whose marker derivatives consistently precede the count derivative with substantial station coverage, yielding actionable lead times from tens to several hundred minutes depending on event morphology. Some descriptors exhibit event dependence: permutation and spectral entropy are strongly anticipatory in the extreme May 2024 episode yet become closer to contemporaneous in the rapid March 2024 case. Heatmaps, violin plots, and station overlays corroborate these patterns and clarify how morphology modulates the available lead. The approach is reproducible from open code, operates on native station units without cross–station homogenization, and remains qualitatively stable under sensitivity sweeps of window, smoothing, and detector parameters. These results support derivative–aligned invariant panels as practical early–warning tools that complement amplitude thresholds and provide a principled basis for nowcasting workflows in space–weather operations.
@@ 1 Introduction
The present work focuses on a set of topological–geometric and complexity measures that are both theoretically grounded and available in robust implementations. We consider Shannon, approximate, sample, permutation, and spectral entropies; Lempel–Ziv complexity; the Hurst exponent and its estimation via detrended fluctuation analysis (DFA); correlation dimension; and fractal dimensions due to Higuchi, Katz, and Petrosian. These measures quantify uncertainty, regularity, memory, roughness, and dimensionality at multiple scales. Comparative and tutorial studies highlight their interpretability and practical diversity, including guidance on parameterization and robustness in noisy, finite data (Unakafova and Keller, 2013; Amigó and Tempesta, 2022; Raubitzek and Neubauer, 2021). Classical nonlinear-dynamics estimators such as the information/correlation dimension provide geometric grounding for state-space structure (Farmer, 1982), while DFA-based scaling analysis links fluctuation growth to long-range dependence in a way that is compatible with weak nonstationarities commonly observed in environmental and space-physics series (Zhou et al., 2014; Stolz and Keller, 2017; Keller and Sinn, 2009).
Within the entropy family, permutation entropy offers a conceptually simple, ordinal-pattern view that is computationally efficient and has been used widely in biomedical and physical time series; extensions to non-uniform embeddings further improve sensitivity in systems with multiple characteristic scales (Unakafova and Keller, 2013; Tao et al., 2018). Shannon and spectral entropies capture distributional and frequency-domain disorder, whereas approximate and sample entropy emphasize conditional irregularity in finite sequences, with well-known trade-offs in parameter choice and data-length requirements (Amigó and Tempesta, 2022; Unakafova and Keller, 2013). Lempel–Ziv complexity provides a symbolization-based measure of algorithmic novelty growth and has proved effective as a complementary descriptor in large feature sets (Fulcher et al., 2013; Lubba et al., 2019). Newer variants, such as SVD entropy, and methodological developments like complexity–entropy maps broaden the diagnostic palette for distinguishing dynamical regimes (Strydom et al., 2021; Ribeiro et al., 2017; Raubitzek and Neubauer, 2021).
@@
Related work in space physics emphasizes the usefulness of complexity-science tools for diagnosing the near-Earth electromagnetic environment and for characterizing transitions across geomagnetic activity levels (Balasis et al., 2023). In cosmic-ray research specifically, multifractal analyses and cross-correlation studies between GCRs and heliospheric parameters have reported persistent correlations and scale-dependent structure consistent with diffusion and convection processes modulated by evolving magnetic turbulence (Sierra-Porta et al., 2024; Sierra-Porta, 2022, 2025). Complementary literature on feature-based time-series learning shows that entropy, fractal, and scaling features can be integrated into supervised and unsupervised workflows, improving detection and ranking of candidate precursors while providing a common ground for cross-station comparisons (Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Raubitzek and Neubauer, 2021). Methodological reviews focused on permutation, approximate, and sample entropy, as well as related complexity measures, provide practical insight into sensitivity to noise, embedding and window choices, and computational efficiency—considerations that are critical when working with neutron-monitor data at varying cadences (Unakafova and Keller, 2013; Amigó and Tempesta, 2022). In parallel, comparative studies of entropy-like descriptors across fields (biomedicine, ecology, finance) illustrate their transferability and the value of multi-measure ensembles, from classical estimators to more recent quantities like bubble and PCA-based entropies (Manis et al., 2021; Panda et al., 2023; Strydom et al., 2021; Ribeiro et al., 2017).
@@
Against this backdrop, our contribution is to develop and evaluate a systematic pipeline that anticipates FDs using temporal markers derived from topological–geometric invariants computed on GCR time series across multiple stations and events. First, we compute sliding-window estimates of entropy-, fractal-, scaling-, and complexity-based measures on each series. Second, we assemble independent marker series for every invariant. Third, we summarize temporal leads in minutes by aligning each marker’s derivative to the derivative-defined onset of the original counts and report the fraction of stations that display sustained pre-onset excursions according to a fixed, robust threshold rule. The novelty lies in operationalizing an early-warning workflow grounded in invariant features—rather than raw-signal thresholds—by comparing entire families of measures within a unified feature framework and prioritizing those markers that most consistently and operationally precede FD onsets across stations and events. This approach yields a principled basis for nowcasting-oriented selection and cross-station generalization, building on the comparative methodology established in feature-based time-series analysis (Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Raubitzek and Neubauer, 2021; Amigó and Tempesta, 2022).
Finally, because FD onsets reflect changes in transport conditions and effective diffusion barriers, they are natural testbeds for precursor discovery. The working hypothesis is that invariant-based markers—capturing changes in complexity, roughness, memory, or dimensionality—will exhibit detectable excursions ahead of amplitude drops when tracked with windows commensurate with interplanetary-disturbance scales. We therefore translate descriptive invariants into operational, station-level predictors of FD onsets via a robust, threshold-and-persistence rule on pre-onset windows, integrating insights from nonlinear time-series analysis, comparative feature learning, and recent space-physics reviews (Balasis et al., 2023; Sierra-Porta, 2022; Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Unakafova and Keller, 2013; Amigó and Tempesta, 2022; Raubitzek and Neubauer, 2021; Zhou et al., 2014; Sinai, 2009; Ribeiro et al., 2017; Manis et al., 2021; Strydom et al., 2021; Panda et al., 2023; Tao et al., 2018; Krakovská and Krakovská, 2016; Hasegawa et al., 2013; Graff et al., 2013).
@@ 2 Data
We use minute–resolution galactic cosmic ray (GCR) measurements from the Neutron Monitor Database (NMDB) for three independent Forbush decrease (FD) episodes: 23 April 2023 (2023-04-23), 24 March 2024 (2024-03-24), and 10 May 2024 (2024-05-10). These cases were chosen to span distinct drivers and morphologies while providing broad multi–station coverage at the native one–minute cadence of NMDB.
For 2023-04-23, a coronal mass ejection (CME) impacted Earth and produced a clear global GCR depression followed by a gradual, multi–day recovery. The 2024-03-24 episode is associated with a fast CME arrival and a G4–class geomagnetic storm; it is notable for an unusually rapid decrease and recovery compared to typical FDs, with pronounced station–dependent anisotropy. The 2024-05-10 event was triggered by the shock of an interplanetary CME (ICME) and ranks among the largest recent FDs (up to about 15.7% at $\sim$10 GV rigidity), coinciding with a severe geomagnetic storm and followed by a ground–level enhancement (GLE 74) on 11 May 2024 during the recovery phase.
The multi–station corpus comprises 20 neutron monitors with usable data for at least one of the target events. Station codes included in this analysis are: AATB, APTY, CALG, DOMB, DOMC, INVK, IRK2, IRK3, JUNG1, KERG, KIEL2, LMKS, MWSN, MXCO, NAIN, NEWK, OULU, PWNK, SOPO, SOPB. Coverage by event, defined as stations with continuous minute–resolution data across the analysis window, is as follows: 10 stations for 2023-04-23, 11 for 2024-03-24, and 11 for 2024-05-10. Full per–station availability and derived marker series are provided in the project repository as supplementary material.
For each FD we extract a symmetric time window around the nominal onset to capture a pre–event baseline, the main decrease, and the early recovery. Timestamps are harmonized to a common one–minute grid per event, and stations showing discontinuities or extended gaps in the immediate pre–onset segment are excluded to avoid bias in marker estimation. Minute–resolution counts are used in their station–specific native units as provided by NMDB (no within–station z–scoring of the count signal). Downstream analyses operate on sliding–window invariant series computed from these counts; each invariant time series is then z–standardized within station prior to taking its first derivative, so that temporal alignment is assessed relative to each station’s own baseline. FD onset per station is defined operationally as the time of the minimum of the first derivative of the smoothed count series, and all leads are reported in minutes (negative values indicate that the marker precedes the onset).
To illustrate the three target FDs, Figure 1 shows minute–resolution GCR count series from representative NM stations near each event date. Vertical dashed lines mark the FD onset window used to align the sliding–window markers. The three panels highlight distinct morphologies: a typical CME–driven drop with gradual recovery (2023–04–23), a fast decrease and unusually rapid rebound concurrent with a G4 storm (2024–03–24), and an extreme case with large amplitude and severe storm context (2024–05–10).
Figure 1: Representative minute–resolution GCR count series around the three FD events analyzed in this work. Vertical dashed lines indicate the FD onset bracket used in this study. Series are NMDB one–minute counts; axes are station–specific native units.
@@ 3 Methods
Let $x_{s}(t)$ denote the minute–resolution GCR count at station $s$ at time index $t$ within an event–specific window that covers pre–onset baseline, the main drop, and early recovery. We retain only contiguous segments with no missing minute–ticks in the immediate pre–onset region. To reduce high–frequency noise while preserving abrupt transitions, we apply an exponentially weighted moving average (EWM) with smoothing parameter $\alpha=0.15$,
$\hat{x}_{s}(t)\;=\;\alpha\,x_{s}(t)\;+\;(1-\alpha)\,\hat{x}_{s}(t-1),$
@@
initialized at the first time point of the event window. All subsequent analyses operate on these EWM-smoothed counts.
To emphasize dynamical change, we work primarily with the derivative of the smoothed signal. With one–minute cadence, we use the central finite difference
$\dot{x}_{s}(t)\;=\;\frac{\hat{x}_{s}(t+1)-\hat{x}_{s}(t-1)}{2},\quad t\in[t_{\mathrm{start}}+1,\,t_{\mathrm{end}}-1],$
using one–sided differences at the endpoints. For robustness to outliers we optionally Winsorize $\dot{x}_{s}(t)$ at the 1st–99th percentiles within the event window; all results reported are stable to this choice.
We detect FD onset at the station level as the time of the minimum of the first derivative of the EWM–smoothed count series, i.e., $t_{0}^{(s)}:=\arg\min_{t}\dot{x}_{s}(t)$, within an event–specific search bracket. Leads for markers are computed per station as time differences relative to $t_{0}^{(s)}$ and are reported in minutes (negative values indicate that the marker precedes the onset).
@@
For each station $s$, window length $W$ (minutes), and stride $\Delta$ (minutes), we compute a family of topological–geometric and complexity invariants on the EWM-smoothed counts. For a window centered at $\tau$ we define
$M_{k}^{(s)}(\tau)\;=\;\mathcal{F}_{k}\!\left(\tilde{x}_{s}\big[\tau-W/2:\tau+W/2\big]\right),\quad\tau\in\mathcal{T}_{W},$
where $k$ indexes the invariant and $\mathcal{T}_{W}$ is the set of window centers that fit in the event window. Unless stated otherwise we set $W=130$ and $\Delta=1$ minute, matching the exploratory figures. We consider the following families, with standard parameterizations (details below):
  • Information–theoretic: Shannon entropy, spectral entropy, approximate entropy, sample entropy, permutation entropy.
  • Algorithmic: Lempel–Ziv complexity.
  • Scaling/memory: Hurst exponent (including DFA estimation (Zhou et al., 2014; Sinai, 2009)).
  • Geometric/fractal: correlation dimension (Farmer, 1982); fractal dimensions of Higuchi, Katz, and Petrosian.
Each $M_{k}^{(s)}(\tau)$ is treated as a marker series. To parallel the treatment of the raw counts, we compute a smoothed first difference
$\dot{M}_{k}^{(s)}(\tau)\;=\;\frac{M_{k}^{(s)}(\tau+\Delta)-M_{k}^{(s)}(\tau-\Delta)}{2\Delta},$
after applying an EWM with the same $\alpha$ to $M_{k}^{(s)}(\tau)$ (defined on the discrete grid $\mathcal{T}_{W}$).
We define station-level onset as the time of the minimum of the first derivative of the EWM-smoothed count series $\dot{x}_{s}(t)$ within an event-specific search bracket. For each invariant $M_{k}$, we first EWM–smooth the windowed series, then z–standardize it within station, and finally take its central finite–difference derivative. The lead for a station–marker pair is the time difference (in minutes) between the minimum of $\dot{M}_{k}$ and the minimum of $\dot{x}_{s}$; negative values mean the marker precedes the onset.
Permutation entropy uses ordinal patterns of embedding dimension $m=3$ and delay $d=1$ (robust alternatives are reported in sensitivity analyses). Sample/approximate entropy use $m=2$ and tolerance $r=0.2\,\sigma$ of the windowed data. Spectral entropy is computed from the normalized periodogram of the window. Lempel–Ziv uses binary symbolization by median thresholding within the window. The Hurst (long-range dependence) properties of the series were estimated using multiple complementary approaches to increase robustness against nonstationarity and intermittency. Primary estimates were obtained via detrended fluctuation analysis (DFA) using polynomial detrending orders 1–2 and linear fits of the fluctuation function on scales within [8,W/2][8,\,W/2] minutes. In addition, we computed classical rescaled-range (R/S) estimates directly derived from Hurst’s original definition and a generalized Hurst exponent $H_{q}$ for a set of moment orders $q$ to probe multifractal scaling. Higuchi’s method uses $k_{\max}=10$. Katz and Petrosian dimensions follow standard definitions on the windowed signal. Correlation dimension uses delay embedding with $(m,d)$ chosen by the first minimum of the auto–mutual information (bounded by $m\leq 8$) and false–nearest–neighbors stopping; scaling is fit on the linear region of $\log C(r)$ vs. $\log r$.
Because FD onsets manifest as rapid negative excursions in the count rate, we examine each marker in the temporal neighborhood of the station–level onset and summarize departures from the pre–event baseline via an operational rule: on a pre–onset window [L,0)[{-}L,0) minutes relative to $t_{0}^{(s)}$, we flag a station–marker pair if the robust z–score of the marker derivative exceeds a fixed threshold (bilateral) for at least $d$ consecutive minutes (cf. parameter choices below). This yields station–level leads in minutes and a per–marker fraction of flagged stations, without relying on cross–correlation or hypothesis tests.
For each invariant $k$ we summarize anticipatory behavior by: (i) the median lead $\widetilde{\ell}_{k}$ (in minutes, negative = precedes) across stations; (ii) the percentage of stations flagged by the operational rule in [L,0)[{-}L,0), denoted $\mathrm{Sig}_{k}[\%]$; and (iii) a robustness score $R_{k}$ that combines coverage, promptness, and concentration:
$R_{k}\;=\;w_{1}\,\mathrm{norm01}\!\big(\mathrm{Sig}_{k}[\%]\big)\;+\;w_{2}\,\mathrm{norm01}\!\big(|\widetilde{\ell}_{k}|\big)\;+\;w_{3}\,\mathrm{norm01}\!\big(-\mathrm{IQR}(\ell)\big),$
with weights $(w_{1},w_{2},w_{3})=(0.5,0.3,0.2)$. Here $\mathrm{norm01}(\cdot)$ denotes a min–max scaling over markers within the summary (a constant input maps to $0.5$ to avoid degenerate scores). Markers are ranked by $R_{k}$; ties are broken by $\mathrm{Sig}_{k}[\%]$ and then by more negative $\widetilde{\ell}_{k}$.
We probe robustness to methodological choices by varying window length $W\in\{90,130,180\}$ minutes; stride $\Delta\in\{1,5\}$ minutes; EWM coefficient $\alpha\in\{0.10,0.15,0.25\}$; derivative scheme (central vs. one–sided); and station–inclusion rules (gap tolerance $\leq 2$ contiguous minutes). We also assess stability of the operational detector by varying the pre–onset window $L\in\{360,480\}$ minutes, the robust threshold $Z_{0}\in\{1.3,1.5\}$, and the persistence requirement $d\in\{5,10\}$ minutes. All qualitative conclusions reported in §4 are stable across these perturbations.
All computations are performed on one–minute NMDB series. Default parameters: $W=130$ min, $\Delta=1$ min, EWM $\alpha=0.15$, derivative winsorization $p=0.01$ (1st–99th percentiles), pre–onset window $L=480$ min, robust threshold $Z_{0}=1.3$, and persistence $d=5$ min. For the invariant implementations we follow standard references for DFA/Hurst (Zhou et al., 2014; Sinai, 2009), correlation dimension (Farmer, 1982), permutation/sample/approximate entropy and related choices (Unakafova and Keller, 2013; Amigó and Tempesta, 2022), and widely used fractal estimators (Higuchi, Katz, Petrosian).
For reproducibility, we summarize below the default configuration of the pipeline and the ranges explored in the sensitivity analysis (See Table 1). These defaults were used for all results unless otherwise noted; variations are explicitly reported in the robustness checks.
Table 1: Default pipeline parameters (with sensitivity ranges used in robustness checks).

### v2

@@ Abstract
We develop a feature-based framework to anticipate Forbush decreases (FDs) in one-minute neutron-monitor records by tracking sliding-window invariants from information theory, scaling, and geometry. For each station we compute marker time series—including Shannon, spectral, approximate and sample entropy; Lempel–Ziv complexity; correlation dimension; and Higuchi and Katz fractal dimensions—smooth them with an exponentially weighted moving average, and analyze their within-station standardized first differences. Timing is referenced to an operational alignment time $t_{0}$ defined as the minimum of the smoothed count first difference, and marker leads are reported in minutes ($\ell^{*}<0$ indicates anticipation). Station-level detectability is defined on a pre-$t_{0}$ window using a robust $z$-score detector with bilateral threshold and persistence, requiring neither cross-correlation nor hypothesis testing.
We apply the pipeline to two FD episodes with broad station coverage (2023-04-23 and 2024-05-10; 28 stations each). Across events, a compact CORE panel exhibits consistently high detection rates and predominantly negative lead distributions, with median leads of order several hours depending on the invariant and event. Lead dispersion across stations is substantial (interquartile ranges typically spanning a few hours), underscoring the value of station-wise criteria and distributional summaries rather than single-station inference. Representative marker trajectories confirm that early flagging corresponds to sustained pre-$t_{0}$ excursions in marker differences, not merely tabulated artifacts.
The approach is reproducible from open code, operates on native station units without cross-station homogenization, and is qualitatively stable to sensitivity sweeps of windowing, smoothing, and detector parameters. These results support derivative-aligned invariant panels as practical early-warning complements to amplitude-threshold methods in space-weather nowcasting.
@@ 1 Introduction
The present work focuses on a set of topological–geometric and complexity measures that are both theoretically grounded and available in robust implementations. We consider Shannon, approximate, sample, permutation, and spectral entropies; Lempel–Ziv complexity; the Hurst exponent and its estimation via detrended fluctuation analysis (DFA); correlation dimension; and fractal dimensions due to Higuchi, Katz, and Petrosian. These measures quantify uncertainty, regularity, memory, roughness, and dimensionality at multiple scales. Comparative and tutorial studies highlight their interpretability and practical diversity, including guidance on parameterization and robustness in noisy, finite data (Unakafova and Keller, 2013; Amigó and Tempesta, 2022; Raubitzek and Neubauer, 2021a). Classical nonlinear-dynamics estimators such as the information/correlation dimension provide geometric grounding for state-space structure (Farmer, 1982), while DFA-based scaling analysis links fluctuation growth to long-range dependence in a way that is compatible with weak nonstationarities commonly observed in environmental and space-physics series (Zhou et al., 2014; Stolz and Keller, 2017; Keller and Sinn, 2009).
Within the entropy family, permutation entropy offers a conceptually simple, ordinal-pattern view that is computationally efficient and has been used widely in biomedical and physical time series; extensions to non-uniform embeddings further improve sensitivity in systems with multiple characteristic scales (Unakafova and Keller, 2013; Tao et al., 2018). Shannon and spectral entropies capture distributional and frequency-domain disorder, whereas approximate and sample entropy emphasize conditional irregularity in finite sequences, with well-known trade-offs in parameter choice and data-length requirements (Amigó and Tempesta, 2022; Unakafova and Keller, 2013). Lempel–Ziv complexity provides a symbolization-based measure of algorithmic novelty growth and has proved effective as a complementary descriptor in large feature sets (Fulcher et al., 2013; Lubba et al., 2019). Newer variants, such as SVD entropy, and methodological developments like complexity–entropy maps broaden the diagnostic palette for distinguishing dynamical regimes (Strydom et al., 2021; Ribeiro et al., 2017; Raubitzek and Neubauer, 2021a).
@@
Related work in space physics emphasizes the usefulness of complexity-science tools for diagnosing the near-Earth electromagnetic environment and for characterizing transitions across geomagnetic activity levels (Balasis et al., 2023). In cosmic-ray research specifically, multifractal analyses and cross-correlation studies between GCRs and heliospheric parameters have reported persistent correlations and scale-dependent structure consistent with diffusion and convection processes modulated by evolving magnetic turbulence (Sierra-Porta et al., 2024; Sierra-Porta, 2022, 2025). Complementary literature on feature-based time-series learning shows that entropy, fractal, and scaling features can be integrated into supervised and unsupervised workflows, improving detection and ranking of candidate precursors while providing a common ground for cross-station comparisons (Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Raubitzek and Neubauer, 2021a). Methodological reviews focused on permutation, approximate, and sample entropy, as well as related complexity measures, provide practical insight into sensitivity to noise, embedding and window choices, and computational efficiency—considerations that are critical when working with neutron-monitor data at varying cadences (Unakafova and Keller, 2013; Amigó and Tempesta, 2022). In parallel, comparative studies of entropy-like descriptors across fields (biomedicine, ecology, finance) illustrate their transferability and the value of multi-measure ensembles, from classical estimators to more recent quantities like bubble and PCA-based entropies (Manis et al., 2021; Panda et al., 2023; Strydom et al., 2021; Ribeiro et al., 2017).
@@
Against this backdrop, our contribution is to develop and evaluate a systematic pipeline that anticipates FDs using temporal markers derived from topological–geometric invariants computed on GCR time series across multiple stations and two FD episodes. We compute sliding-window estimates of entropy-, fractal-, and complexity-based measures, form independent marker series, and quantify station-wise temporal leads in minutes by aligning marker derivatives to a derivative-defined reference time of the original counts. We further report the fraction of stations that exhibit sustained pre-onset excursions under a fixed robust threshold-and-persistence rule, enabling operational comparison and selection of a compact marker panel within a unified feature framework (Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Raubitzek and Neubauer, 2021a; Amigó and Tempesta, 2022; Balasis et al., 2023; Gil et al., 2021).
Finally, because FD onsets reflect changes in transport conditions and effective diffusion barriers, they are natural testbeds for precursor discovery. The working hypothesis is that invariant-based markers—capturing changes in complexity, roughness, memory, or dimensionality—will exhibit detectable excursions ahead of amplitude drops when tracked with windows commensurate with interplanetary-disturbance scales. We therefore translate descriptive invariants into operational, station-level predictors of FD onsets via a robust, threshold-and-persistence rule on pre-onset windows, integrating insights from nonlinear time-series analysis, comparative feature learning, and recent space-physics reviews (Balasis et al., 2023; Sierra-Porta, 2022; Fulcher et al., 2013; Fulcher and Jones, 2014; Lubba et al., 2019; Unakafova and Keller, 2013; Amigó and Tempesta, 2022; Raubitzek and Neubauer, 2021b; Zhou et al., 2014; Sinai, 2009; Ribeiro et al., 2017; Manis et al., 2021; Strydom et al., 2021; Panda et al., 2023; Tao et al., 2018; Krakovská and Krakovská, 2016; Hasegawa et al., 2013; Graff et al., 2013).
@@ 2 Data
We use 1-minute galactic cosmic ray (GCR) count rates from the Neutron Monitor Database (NMDB) for two Forbush decrease (FD) episodes: 23 April 2023 (2023-04-23) and 10 May 2024 (2024-05-10). These two cases were selected to represent distinct FD morphologies while retaining broad multi-station coverage at the native 1-minute cadence.
For 2023-04-23, a coronal mass ejection (CME) impacted Earth and produced a clear global GCR depression followed by a gradual, multi–day recovery. The 2024-05-10 event was triggered by the shock of an interplanetary CME (ICME) and ranks among the largest recent FDs (up to about 15.7% at $\sim$10 GV rigidity), coinciding with a severe geomagnetic storm and followed by a ground–level enhancement (GLE 74) on 11 May 2024 during the recovery phase (Abunina et al., 2025; Papaioannou et al., 2025).
For each event we analyze a curated subset of $N_{s}=28$ stations with continuous 1-minute coverage across the analysis window (see Methods for the continuity criteria). Minute-resolution counts are used in their station-specific native units as provided by NMDB. Downstream analyses operate on sliding-window invariant series computed from these counts; each invariant series is then standardized within station prior to derivative-based lead extraction, so that timing is assessed relative to each station’s own baseline. For each event, the complete set of stations we used for the analysis is: 20230423: [ATHN, DOMB, SOPO, SOPB, THUL, NAIN, INVK, FSMT, TXBY, NRLK, APTY, OULU, CALG, DOMC, YKTK, KERG, NEWK, MXCO, NANM, KIEL2, BKSN, JUNG, TERA, LMKS, IRK2, IRKT, JUNG1, PWNK]; 20240510: [MXCO, DOMB, SOPO, SOPB, MWSN, MWSB, THUL, PWNK, NAIN, TXBY, APTY, OULU, DOMC, KERG, CALG, KIEL2, NANM, CALM, BKSN, JUNG, JUNG1, YKTK, LMKS, NEWK, TERA, IRK3, DRBS, INVK]. The short names of the stations can be found in https://www.nmdb.eu/nest/ y https://www.nmdb.eu/nest/help.php#helpstations.
Figure 1 shows representative 1-minute count-rate series for each event (THUL for 2023-04-23 and CALG for 2024-05-10), highlighting the contrasting morphologies that motivate a station-wise, derivative-aligned alignment time.
Figure 1: Representative 1-minute NMDB count-rate series for the two analyzed FD events: THUL (2023-04-23) and CALG (2024-05-10). The panels illustrate distinct FD morphologies (a moderate event with gradual recovery versus an extreme event with large amplitude and a sharper main decrease), motivating a station-wise, derivative-aligned definition of the reference time used throughout the analysis.
@@ 3 Methods
3.1 Data, smoothing, and operational alignment
Let $x_{s}(t)$ denote the one-minute neutron-monitor count-rate series at station $s$ over an event-specific analysis window covering pre-event baseline, the main decrease, and early recovery. We retain stations with sufficient coverage in the immediate pre-onset interval and treat missing values conservatively by requiring a minimum valid fraction per analysis window (see sensitivity ranges below).
To reduce high-frequency noise while preserving abrupt transitions, we apply an exponentially weighted moving average (EWM) with coefficient $\alpha$,
$\hat{x}_{s}(t)\;=\;\alpha\,x_{s}(t)\;+\;(1-\alpha)\,\hat{x}_{s}(t-1),$ (1)
@@ 3.1 Data, smoothing, and operational alignment
initialized at the first time point of the event window. All subsequent computations operate on $\hat{x}_{s}(t)$.
Rather than adopting a threshold-based physical FD onset definition, we define a station-wise operational alignment time $t_{0}^{(s)}$ anchored in the count-rate dynamics. Specifically, we compute the EWM-smoothed first difference
$\Delta\hat{x}_{s}(t)\;=\;\hat{x}_{s}(t)-\hat{x}_{s}(t-1),$ (2)
and set
$t_{0}^{(s)}\;:=\;\arg\min_{t\in\mathcal{B}}\Delta\hat{x}_{s}(t),$ (3)
where $\mathcal{B}$ is an event-specific search bracket chosen to include the main decrease phase. Throughout the manuscript, $t_{0}^{(s)}$ is used as an alignment reference to synchronize heterogeneous station profiles; it is not claimed to be a universal physical onset time.
3.2 Sliding-window invariants (marker series)
@@
For each station $s$, we compute a set of sliding-window invariants on the smoothed counts $\hat{x}_{s}(t)$ using a causal window of length $W$ minutes evaluated every $\Delta t$ minutes. For evaluation times $t_{j}$ on the grid $\mathcal{T}$, we define the causal window
$\mathcal{W}(t_{j})\;=\;\{\hat{x}_{s}(u):u\in[t_{j}-W,\,t_{j}]\}.$
The marker (invariant) series for metric $k$ is then
$M_{k}^{(s)}(t_{j})\;=\;\mathcal{F}_{k}\!\big(\mathcal{W}(t_{j})\big),\qquad t_{j}\in\mathcal{T},$ (4)
where $\mathcal{F}_{k}$ denotes the estimator for invariant $k$.
This causal construction avoids look-ahead and makes the marker time series compatible with online early-flagging. We initially compute a broader screening set spanning information-theoretic, geometric/fractal, and complexity/embedding families; the main results focus on a compact CORE panel selected for high detection coverage and stable lead distributions across events.
3.3 Marker dynamics and lead definition
To summarize dynamical changes in markers, we smooth each marker series with an EWM (same $\alpha$ unless noted) and compute its first difference on the evaluation grid,
$\Delta M_{k}^{(s)}(t_{j})\;=\;\tilde{M}_{k}^{(s)}(t_{j})-\tilde{M}_{k}^{(s)}(t_{j-1}),$ (5)
where $\tilde{M}_{k}^{(s)}$ is the EWM-smoothed marker series. We define the lead for a station–marker pair as
$\ell_{k,s}^{*}\;=\;t_{k,s}-t_{0}^{(s)},$ (6)
where $t_{k,s}$ is the time of the most negative excursion of $\Delta M_{k}^{(s)}(t)$ within the pre-onset search window $t_{0}^{(s)}-L,\;t_{0}^{(s)})$. Leads are reported in minutes; $\ell_{k,s}^{*}<0$ indicates anticipation (marker change precedes $t_{0}^{(s)}$).
3.4 Operational detector and station-wise coverage
In addition to lead timing, we quantify whether a marker exhibits a sustained pre-$t_{0}$ excursion via an operational detector on the pre-onset window $t_{0}^{(s)}-L,\;t_{0}^{(s)})$. We compute a robust $z$-score of the marker difference series using the median and MAD on the same pre-onset interval,
$z_{k,s}(t)\;=\;\frac{\Delta M_{k}^{(s)}(t)-\mathrm{median}(\Delta M_{k}^{(s)})}{1.4826\,\mathrm{MAD}(\Delta M_{k}^{(s)})},$
and flag a station–marker pair if $|z_{k,s}(t)|\geq Z_{0}$ for at least $d$ consecutive evaluation steps (bilateral threshold with persistence). This yields a station-wise detection indicator and, for each invariant $k$, the detection coverage $\mathrm{Detect}_{k}[\%]$ across the station network.
3.5 Invariant estimators and default parameters
The CORE panel used in §4 includes: Shannon entropy, spectral entropy, sample entropy, approximate entropy, Lempel–Ziv complexity, Higuchi fractal dimension, Katz fractal dimension, and correlation dimension. We follow standard definitions and widely used implementations for entropy-type measures and related parameterizations (Unakafova and Keller, 2013; Amigó and Tempesta, 2022), DFA/Hurst and scaling estimators where used in screening (Zhou et al., 2014; Sinai, 2009), and correlation dimension (Farmer, 1982). Unless otherwise noted: spectral entropy uses a Welch PSD with sampling frequency $sf=1$ (one sample per minute) and normalization; Higuchi uses $k_{\max}=10$; and Lempel–Ziv uses binary symbolization by median thresholding within each window.
3.6 Sensitivity analysis
We probe robustness to methodological choices by varying: window length $W\in\{90,120,180\}$ min; stride $\Delta t\in\{1,5\}$ min; EWM coefficient $\alpha\in\{0.10,0.15,0.25\}$; pre-onset window $L\in\{360,480\}$ min; robust threshold $Z_{0}\in\{1.3,1.5\}$; and persistence $d\in\{3,5,10\}$ evaluation steps. All qualitative conclusions reported in §4 are stable across these perturbations. For clarity and reproducibility, Table 1 summarizes the default pipeline parameters used to produce the results in §4, together with the ranges explored in the sensitivity analysis.
Table 1: Default pipeline parameters (with sensitivity ranges). The marker windows are causal (no look-ahead) and evaluated every $\Delta t$ minutes.